# Number Cookbook: Number Understanding of Language Models and How to Improve It

### v2

@@ Abstract
Large language models (LLMs) can solve an increasing number of complex reasoning tasks while making surprising mistakes in basic numerical understanding and processing (such as $9.11>9.9$). The latter ability is essential for tackling complex arithmetic and mathematical problems and serves as a foundation for most reasoning tasks, but previous work paid little attention to it or only discussed several restricted tasks (like integer addition). In this paper, we comprehensively investigate the numerical understanding and processing ability (NUPA) of LLMs. Firstly, we introduce a benchmark covering four common numerical representations and 17 distinct numerical tasks in four major categories, resulting in 41 meaningful combinations in total. These tasks are derived from primary and secondary education curricula, encompassing nearly all everyday numerical understanding and processing scenarios, and the rules of these tasks are very simple and clear. Through the benchmark, we find that current LLMs fail frequently in many of the tasks. To study the problem, we train small models with existing and potential techniques for enhancing NUPA (such as tokenizers, PEs, and number formats), comprehensively evaluating their effectiveness using our testbed. We also finetune practical-scale LLMs on our proposed NUPA tasks and find that 1) naive finetuning can improve NUPA a lot on many but not all tasks, and 2) surprisingly, techniques designed to enhance NUPA prove ineffective for finetuning pretrained models. We further explore the impact of chain-of-thought techniques on NUPA. Our work provides a more detailed and comprehensive understanding of NUPA in LLMs. Our benchmark and code are released at https://github.com/GraphPKU/number_cookbook.
@@ 1 Introduction
The mathematical and reasoning abilities of large language models (LLMs) are currently quite impressive (OpenAI, 2023; Meta, 2024a; OpenAI, 2024a; Yang et al., 2024a), capable of solving problems at the level of a graduate student or even more difficult ones like olympiad-level problems (He et al., 2024), GAOKAO (a nationwide examination of high school students applying to universities in China) (Zhang et al., 2024b) and college mathematics (Tang et al., 2024). However, upon closer examination of the models’ outputs, we found that although the models demonstrate remarkable proficiency in problem-solving approaches, they often struggle with basic numerical understanding and processing — like a careless student who claims, “I know how to do it, but I didn’t get it right.” Some of these errors are quite surprising, such as believing that $9.11>9.9$ or making mistakes in simple addition $8/7+3/5$. These errors are a major cause of hallucinations when dealing with math, reasoning, and data analysis tasks, as the model presents seemingly correct problem-solving approaches, but ultimately produces incorrect results (Huang et al., 2024; Li et al., 2024b; Jiang et al., 2024b). Therefore, investigating and improving the fundamental “numerical understanding and processing abilities” (NUPA) of models is crucial.
@@
However, in current research, reasoning ability and NUPA are often tested together, both on classic datasets such as GSM8k (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), MMLU (Hendrycks et al., 2020), and in more challenging tests mentioned above. For example, a problem in GSM8k is: “Natalia sold clips to 48 of her friends in April, and then she sold half as many clips in May. How many clips did Natalia sell altogether in April and May?” Solving this problem requires two aspects: on the one hand, mathematical reasoning including understanding the text, extracting relevant information, formulating mathematical equations (or finding other solution methods), solving the equations or executing an algorithm, and obtaining the result; on the other hand, it also requires understanding and processing the numbers provided in the problem or produced as intermediate results at each step, like $48/2=24$ and $48+24=72$. While these two abilities are both essential to correctly solving the problems, tests on such datasets do not distinguish between them.
@@
A more severe issue is that the numerical content is often deliberately simplified in these datasets. In various exam questions (like in the American Invitational Mathematics Examination (Li et al., 2024a)), to focus on assessing students’ understanding of mathematical concepts — such as how to set up the correct equations and apply the right theorems — the numbers in both the questions and answers are often specially chosen to be integers. However, this is not the case in real-world scenarios (Chen et al., 2022a).
@@
Despite the importance of NUPA, there is still a lack of accurate, detailed, and comprehensive formalization, measurement, and analysis of this fundamental capability. In this paper, we take the preliminary step towards formalizing the NUPA of LLMs. We categorize the numerical concepts and operations commonly taught in primary and secondary education into four types of numerical representations: integers, floating-point numbers (finite decimals), fractions, and scientific notation, along with four ability categories comprising 17 tasks. Pairing these representations results in 41 meaningful tasks, forming our NUPA benchmark (Table 1). These representations and tasks cover the most common scenarios involving number understanding and processing, which are typically not challenging for humans, as we read, use, or process such numbers nearly everyday.
@@
On this benchmark, we rigorously test several state-of-the-art LLMs containing GPT-4o (OpenAI, 2024a), Llama-3.1 (Meta, 2024a) and Qwen2 (Qwen Team, 2024). We ask models to directly output the answers without calling external tools. Although the latest LLMs perform well on some easiest tasks, their performance declines significantly as tasks become slightly more complex (such as multiplication, modulo operations, or digit-based calculations), or as the representation of numbers extends beyond basic integers. See Figure 2 of Section 2.4. The overall unsatisfactory performance highlights a pronounced mismatch between the claimed strong mathematical reasoning abilities and the poor practical, everyday numerical understanding and processing abilities of today’s LLMs.
@@
To address this issue, we explore three categories of approaches to enhance the NUPA of models. The first category of techniques aim at improving models’ NUPA during the pretraining stage, including alternative tokenization, specially designed positional encoding (PE) (Haviv et al., 2022; Kazemnejad et al., 2023b; Zhou et al., 2024), changing number formats (like zero-padding, index-hint (Zhou et al., 2023) and reverse representation (Lee et al., 2023; Zhou et al., 2024)). We evaluate and analyze them on our newly introduced benchmark, verifying their effectiveness/ineffectiveness on respective tasks/representations, which extends over previous evaluation mainly on the integer addition/multiplication tasks. Further, we summarize these techniques into three mechanisms: simplifying the reasoning process, aiding digit alignment, and providing regularization, and discuss the potential of these mechanisms to be applied across a broader range of numerical representations.
@@
The second category of approaches aim to improve NUPA for an already trained model. We find that while simple direct finetuning can significantly enhance NUPA performance, applying the aforementioned techniques (PEs, data formats and tokenizers) at this stage may have adverse effects. We test various settings and fine-tuning configurations, but none are able to achieve performance equal to or better than the original model. Our results suggest that these modifications can significantly disrupt the models’ established behavior or conflict with its pre-existing knowledge, leading to a decrease in performance.
Finally, we discuss the potential of using chain-of-thought (CoT) techniques (Wei et al., 2023) for numerical processing. Although CoT methods can break down complex problems into simpler sub-tasks and significantly increase the likelihood of obtaining correct answers, their drawbacks — such as consuming a large context window and requiring extended processing time — become particularly apparent in numerical tasks. We test a general CoT method known as RFFT (Hu et al., 2024), and find that for more complex tasks (such as multiplication and fraction addition), chain-of-thought methods face scalability challenges, making them difficult to be applied in practical scenarios. It is noteworthy that in this paper, we do not discuss tool use methods (Schick et al., 2023; Lu et al., 2023a) for NUPA as 1) we want to study the self-contained NUPA of LLMs, 2) calling external tools whenever encountering numbers increases the inference latency (Xu et al., 2024), and 3) we believe NUPA without tools is a necessary ability of AGI.
@@ 2.1 Number representation
As we discuss above, we believe that the educational curricula on the (Chinese) primary and secondary school levels is a valuable reference for determining the essential NUPAs that LLMs should master. We identify four number formats in these curricula that are both common and sufficient to cover most practical scenarios.
@@ 2.2 Tasks in four ability categories
Another aspect of NUPA is defining the tasks that the model needs to handle. The tasks should have clear calculation rules — Because these tasks are extracted from the education curricula, students who have completed the stage of education are expected to solve them. Furthermore, most practical numerical processing tasks should either fall within these tasks or can be easily transformed into some of them. We propose 17 tasks across four ability categories. The complete task list is shown in Table 1 and we provide a more detailed discussion in Appendix A.1.2 an example for each task in Appendix A.1.3. Below we discuss the rationales for including some tasks in detail.
@@
The combination of representations and tasks ultimately result in a total of 41 meaningful pairs. Without confusion, we also call them tasks. The tasks receive either one or two numbers as inputs and return a number as result and the input numbers and results share the same representation for most tasks unless otherwise stated (refer to Appendix A.1.4). The remaining combinations are excluded due to being excessively complex, uncommon, inapplicable, or redundant with other tasks. For further details, see the discussion in Appendix A.1.5.
The difficulty of each task depends not only on the nature of the task itself but also on the length of the numbers to be processed — longer tasks involve more steps of internal operations, as well as longer inputs and outputs. Therefore, we test on different problem lengths. For tasks that are inherently more difficult, we limit the size of the problem to 1-20 digits, and for easier tasks to 1-100 digits. (For which tasks are considered difficult or easy, please refer to the Appendix A.1.6.)
We generated 1,000 questions for each task and each length. Unlike some previous works that set the lengths of two numbers to be the same, in our tests, the length $L$ of a question is determined by the longer of the two numbers, while the length of the shorter number follows a uniform distribution between $1$ and $L$. We implemented additional handling to ensure that generated problems do not result in overly simple, complex, or meaningless results. Some tasks are further split into a hard and an easy version. More details about generating the benchmark are provided in Appendix A.1.7.
@@ 2.4 Performance of current LLMs
We test some commonly used LLMs on our benchmark, including three Llama models: Llama-2-7b, Llama-3.1-8b and Llama-3.1-70b (Meta, 2024a), one of the most popular open-source model families from Meta; Mixtral-8$\times$7B (Jiang et al., 2024a), a strong MoE model; and Qwen2-2B and Qwen-72B (Qwen Team, 2024) which are also open-source models that are believed to have strong math abilities. Finally, we also test state-of-the-art commercial models GPT-4o-2024-08-06 and GPT-4o-mini-2024-07-18 (OpenAI, 2024a). We use prompts to control models to directly output result numbers without relying on external tools or CoT. The prompts used for each model and task are included in Appendix A.2. We select the results of some typical tasks in each category in Figure 2, while the complete results and discussion on all metrics are shown in Appendix A.3. And we also provide an interactive performance page where readers can choose tasks they are interested in. Here, we mainly focus on the zero-shot performance while we discuss about few-shot performance in Appendix A.3.1. We have several observations regarding the results:
Figure 2: Parts of performance of state-of-the-art LLMs on NUPA benchmark. “-ft” denotes a Llama model we finetuned on these tasks. (See Section 3.4)
The best model performs well on typical tasks, but its performance declines on more specialized tasks. We find that GPT-4o, GPT-4o-mini and Qwen2 handle typical tasks, such as integer addition, float addition, integer max, and integer length, with high accuracy in the S and M ranges. This aligns with their strong performance on various mathematical datasets. However, their accuracy drops sharply when working with less common representations, like fractions and scientific notation, with average accuracy falling below 20%, even for the shortest S-range (1-4 digits). Similarly, for tasks such as significant figures, modulo operations, and digit-based calculations, their performance was unsatisfactory. This highlights the current limitations of LLMs in understanding numerical diversity and complexity. Despite their good performance on a narrow set of numerical tasks, they struggle with many others, failing to produce accurate results in these areas.
@@
Length impedes learning both individual digits and overall length. To understand why models struggle with longer input numbers, we examine digit match and dlength performance in Figure 6 and Figure 7 in Appendix A.3. These metrics reveal that length affects both the accuracy of individual digits (digit match) and the answer’s overall length (dlength), with variations across tasks. For example, GPT-4o and Llama-3.1 display consistently low dlength in the add-integer task, with digit match decreasing sharply as length increases, suggesting that length primarily impacts per-digit accuracy on this task. Conversely, in the max-float task, dlength increases significantly with length (about 30-60 in the XL range), while digit match has still 60% in the XL range. Note that since missing digits are treated as errors, this 0.6 digit match is likely due to these missing digits. This suggests that the main challenge here lies in generating answers of the correct length, rather than individual digit accuracy. In other tasks like fraction, both length and digit accuracy issues arise, as reflected in rising dlength and declining digit match.
@@
“Digit” is more challenging than expected. We were surprised to find that LLMs struggle to fully grasp “digits”. For instance, in the “get digit” task, where the model is asked to return the $i$-th digit of a long integer, performance drops significantly as the length of the number increases. This suggests that current LLMs lack a consistent ability to just find a digit. Note that the performance is good in the shorter S-range, which indicates that the models can at least comprehend the task instruction. In the XL-range, GPT-4o achieves only 20% accuracy, barely above the random guessing 10% baseline (since the correct answer is always a digit between 0 and 9). This fundamental limitation may explain why current LLMs struggle with numerical understanding and processing, especially as task complexity and input length increase. If a model cannot reliably identify a specific digit in a given number, it casts doubt on its ability to generalize to more complex arithmetic tasks, such as addition.
@@
We also have some interesting observations: (1) LLMs find the “max-hard” task easier than “max” with integer inputs. The difference between these tasks is that in the max task, the two numbers often differ in length, whereas in max-hard, they are always the same length and share some left-most digits, requiring more digits to be compared. While max-hard intuitively seems more difficult, models actually perform better on it. This is likely because they struggle to effectively use sequence length information, as reflected in their weaker performance on the “length” tasks in the longer ranges. It suggests that models might process tasks in different ways from humans. They could have to compare two numbers digit by digit. In this situation, the “harder” subtasks are actually easier because the numbers are already aligned. (2) GPT-4o and GPT-4o-mini show nearly identical performance across most tasks, similar to the comparison between Qwen2-72B and Qwen2-7B. This suggests that once a model reaches a certain size, NUPA performance relies more on factors like architecture design, training strategies, data diversity, and post-training refinements, rather than simply on increasing model size.

### v3

@@ Abstract
Large language models (LLMs) can solve an increasing number of complex reasoning tasks while making surprising mistakes in basic numerical understanding and processing (such as $9.11>9.9$). The latter ability is essential for tackling complex arithmetic and mathematical problems and serves as a foundation for most reasoning tasks, but previous work paid little attention to it or only discussed several restricted tasks (like integer addition). In this paper, we comprehensively investigate the numerical understanding and processing ability (NUPA) of LLMs. Firstly, we introduce a benchmark covering four common numerical representations and 17 distinct numerical tasks in four major categories, resulting in 41 meaningful combinations in total. These tasks are derived from primary and secondary education curricula, encompassing nearly all everyday numerical understanding and processing scenarios, and the rules of these tasks are very simple and clear. Through the benchmark, we find that current LLMs fail frequently in many of the tasks. To study the problem, we train small models with existing and potential techniques for enhancing NUPA (such as tokenizers, positional encoding, and number formats), comprehensively evaluating their effectiveness using our testbed. We also finetune practical-scale LLMs on our proposed NUPA tasks and find that 1) naive finetuning can significantly improve NUPA on many but not all tasks, and 2) surprisingly, techniques designed to enhance NUPA prove ineffective for finetuning pretrained models. We further explore the impact of chain-of-thought techniques on NUPA. Our work provides a more detailed and comprehensive understanding of NUPA in LLMs. Our benchmark and codes are released at https://github.com/GraphPKU/number_cookbook.
@@ 1 Introduction
The mathematical and reasoning abilities of large language models (LLMs) are currently quite impressive (OpenAI, 2023; Meta, 2024a; OpenAI, 2024a; Yang et al., 2024a), capable of solving problems at the level of a high-school student or even more difficult ones like GAOKAO (a nationwide examination of high school students applying to universities in China) (Zhang et al., 2024b), Olympiad-level problems (He et al., 2024), and college mathematics (Tang et al., 2024). However, upon closer examination of the models’ outputs, we find that although the models demonstrate remarkable proficiency in problem-solving approaches, they often struggle with basic numerical understanding and processing — like a careless student who claims, “I know how to do it, but I didn’t get it right.” Some of these errors are quite surprising, such as believing that $9.11>9.9$ or making mistakes in simple addition $8/7+3/5$. These errors are a major cause of hallucinations when dealing with math, reasoning, and data analysis tasks, as the model presents seemingly correct problem-solving approaches, but ultimately produces incorrect results (Huang et al., 2024; Li et al., 2024b; Jiang et al., 2024). Therefore, investigating and improving the fundamental “numerical understanding and processing abilities” (NUPA) of models is crucial.
@@
However, in current research, reasoning ability and NUPA are often tested together, both on classic datasets such as GSM8k (Cobbe et al., 2021), MATH (Hendrycks et al., 2021b), MMLU (Hendrycks et al., 2021a), and in more challenging tests mentioned above. For example, a problem in GSM8k is: “Natalia sold clips to 48 of her friends in April, and then she sold half as many clips in May. How many clips did Natalia sell altogether in April and May?” Solving this problem requires two aspects: on the one hand, mathematical reasoning including understanding the text, extracting relevant information, formulating mathematical equations (or finding other solution methods), solving the equations or executing an algorithm, and obtaining the result; on the other hand, it also requires understanding and processing the numbers provided in the problem or produced as intermediate results at each step, like $48/2=24$ and $48+24=72$. While these two abilities are both essential to correctly solving the problems, tests on such datasets do not distinguish between them.
@@
A more severe issue is that the numerical content is often deliberately simplified in these datasets. In various exam questions (like in the American Invitational Mathematics Examination (Li et al., 2024a)), to focus on assessing students’ understanding of mathematical concepts — such as how to set up the correct equations and apply the right theorems — the numbers in both the questions and answers are often specially chosen to be integers. However, this is not the case in real-world scenarios (Chen et al., 2021).
@@
Despite the importance of NUPA, there is still a lack of accurate, detailed, and comprehensive formalization, measurement, and analysis of this fundamental capability. In this paper, we take the preliminary step towards formalizing the NUPA of LLMs. We categorize the numerical concepts and operations from primary and secondary education into four representations: integers, floating-point numbers (finite decimals), fractions, and scientific notation, along with four ability categories comprising 17 tasks. Pairing these representations results in 41 meaningful tasks, forming our NUPA benchmark (Table 1). These representations and tasks cover the most common scenarios involving number understanding and processing, which are typically not challenging for humans, as we read, use, or process such numbers nearly every day.
@@
On this benchmark, we rigorously test several state-of-the-art LLMs containing GPT-4o (OpenAI, 2024a), Llama-3.1 (Meta, 2024a) and Qwen2 (Qwen Team, 2024). We ask the models to directly output the answers without calling external tools. Although the latest LLMs perform well on some of the easiest tasks, their performance declines significantly as tasks become slightly more complex (such as multiplication, modulus operations, or digit-based calculations), or as the representation of numbers extends beyond basic integers. See Figure 2 of Section 2.4. The overall unsatisfactory performance highlights a pronounced mismatch between the claimed strong mathematical reasoning abilities and the poor practical, everyday numerical understanding and processing abilities of today’s LLMs.
@@
To address this issue, we explore three categories of approaches to enhance the NUPA of models. The first category of techniques aims at improving models’ NUPA during the pretraining stage, including alternative tokenization, specially designed positional encoding (PE) (Haviv et al., 2022; Kazemnejad et al., 2023; Zhou et al., 2024b), changing number formats (like zero-padding, index-hint (Zhou et al., 2024a) and reverse representation (Lee et al., 2024; Zhou et al., 2024b)). We evaluate and analyze them on our newly introduced benchmark, verifying their effectiveness/ineffectiveness on respective tasks/representations, which extends over previous evaluation mainly on the integer addition/multiplication tasks. Further, we summarize these techniques into three mechanisms: simplifying the reasoning process, aiding digit alignment, and providing regularization, and discuss the potential of these mechanisms to be applied across a broader range of numerical representations.
@@
The second category of approaches aim to improve NUPA for an already trained model. We find that while simple direct finetuning can significantly enhance NUPA performance, applying the aforementioned techniques (PEs, data formats and tokenizers) at this stage may have adverse effects. We test various settings and finetuning configurations, but none are able to achieve performance equal to or better than the original model. Our results suggest that these modifications can significantly disrupt the models’ established behavior or conflict with its pre-existing knowledge, leading to a decrease in performance.
Finally, we discuss the potential of using chain-of-thought (CoT) techniques (Wei et al., 2022) for numerical processing. Although CoT methods can break down complex problems into simpler sub-tasks and significantly increase the likelihood of obtaining correct answers, their drawbacks — such as consuming a large context window and requiring extended processing time — become particularly apparent in numerical tasks. We test a general CoT method known as RFFT (Hu et al., 2024), and find that for more complex tasks (such as multiplication and fraction addition), chain-of-thought methods face scalability challenges, making them difficult to be applied in practical scenarios. It is noteworthy that in this paper, we do not discuss tool use methods (Schick et al., 2023; Lu et al., 2023a) for NUPA as 1) we want to study the self-contained NUPA of LLMs, 2) calling external tools whenever encountering numbers increases the inference latency (Xu et al., 2024), and 3) we believe NUPA without tools is a necessary ability of AGI.
@@ 2.1 Number representation
As discussed above, we believe that the educational curricula on the (Chinese) primary and secondary school levels serve as a valuable reference for determining the essential NUPAs that LLMs should master. We identify four number formats in these curricula that are both common and sufficient to cover most practical scenarios.
@@ 2.2 Tasks in four ability categories
Another aspect of NUPA is defining the tasks that models need to handle. The tasks should have clear calculation rules. Furthermore, most practical numerical processing tasks should either fall within these tasks or can be easily transformed into some of them. Extracted from the primary and secondary education curricula, we propose 17 tasks in four ability categories and students who have completed the stage of education are expected to solve them. The complete task list is shown in Table 1 and we provide a more detailed discussion in Appendix A.1.2 and an example for each task in Appendix A.1.3. Below we discuss the rationales for including some tasks in detail.
@@
The combination of representations and tasks ultimately results in a total of 41 meaningful pairs. Without confusion, we refer to each combination as a task. The tasks receive either one or two numbers as inputs and return a number as result, and the input numbers and results share the same representation for most tasks unless otherwise stated (refer to Appendix A.1.4). The remaining combinations are excluded due to being excessively complex, uncommon, inapplicable, or redundant with other tasks. For further details, see the discussion in Appendix A.1.5.
The difficulty of each task depends not only on the nature of the task itself but also on the length of the numbers to be processed — longer tasks involve longer inputs and outputs as well as more steps of internal operations. Therefore, we test on different problem lengths. For tasks that are inherently more difficult, we limit the size of the problem to 1-20 digits, and for easier tasks to 1-100 digits. (For which tasks are considered difficult or easy, please refer to the Appendix A.1.6.)
We generated 1,000 questions for each task and each length. Unlike some previous works that set the lengths of two numbers to be the same, in our tests, the length $L$ of a question is determined by the longer of the two numbers, while the length of the shorter number follows a uniform distribution between $L/2$ and $L$. We implemented additional handling to ensure that generated problems do not result in overly simple, complex, or meaningless results. Some tasks are further split into a hard and an easy version. More details about generating the benchmark are provided in Appendix A.1.7.
@@ 2.4 Performance of current LLMs
We test some commonly used LLMs on our benchmark, including three Llama models: Llama-2-7b, Llama-3.1-8b and Llama-3.1-70b (Meta, 2024a), one of the most popular open-source model families from Meta; Mixtral-8$\times$7B (MistralAI, 2024), a strong MoE model; and Qwen2-7B and Qwen2-72B (Qwen Team, 2024) which are also open-source models that are believed to have strong math abilities. Finally, we also test state-of-the-art commercial models GPT-4o-2024-08-06 and GPT-4o-mini-2024-07-18 (OpenAI, 2024a). We use prompts to control models to directly output result numbers without relying on external tools or CoT. The prompts used for each model and task are included in Appendix A.2. We select the results of some typical tasks in each category in Figure 2, while the complete results [^1] An interactive performance report is shown in https://huggingface.co/spaces/kangshijia/NUPA-Performance. and discussion on all metrics are shown in Appendix A.3. Here, we mainly focus on the zero-shot performance while we discuss few-shot performance in Appendix A.3.1. We have several observations regarding the results:
Figure 2: Parts of performance of state-of-the-art LLMs on NUPA benchmark.1 “-ft” denotes a Llama model we finetuned on these tasks. (See Section 3.4)
The best model performs well on typical tasks, but its performance declines on more specialized tasks. We find that GPT-4o, GPT-4o-mini and Qwen2 handle typical tasks, such as integer addition, float addition, integer max, and integer length, with high accuracy in the S and M ranges. This aligns with their strong performance on various mathematical datasets. However, their accuracy drops sharply when working with less common representations, like fractions and scientific notation, with average accuracy falling below 20%, even for the shortest S-range (1-4 digits). Similarly, for tasks such as significant figures, modulus operations, and digit-based calculations, their performance was unsatisfactory. This highlights the current limitations of LLMs in understanding numerical diversity and complexity. Despite their good performance on a narrow set of numerical tasks, they struggle with many others, failing to produce accurate results in these areas.
@@
Length impedes learning both individual digits and overall length. To understand why models struggle with longer input numbers, we examine digit match and dlength performance in Figure 6 and Figure 7 in Appendix A.3. These metrics reveal that length affects both the accuracy of individual digits (digit match) and the answer’s overall length (dlength), with variations across tasks. For example, GPT-4o and Llama-3.1 display consistently low dlength in the add-integer task, with digit match decreasing sharply as length increases, suggesting that length primarily impacts per-digit accuracy on this task. Conversely, in the max-float task, dlength increases significantly with length (about 30-60 in the XL range), while digit match remains at 60% in the XL range. Note that since missing digits are treated as errors, this 0.6 digit match is likely due to these missing digits. This suggests that the main challenge here lies in generating answers of the correct length, rather than individual digit accuracy. In other tasks like fraction, both length and digit accuracy issues arise, as reflected in rising dlength and declining digit match.
@@
“Digit” is more challenging than expected. We were surprised to find that LLMs struggle to fully grasp “digits”. For instance, in the “get digit” task, where the model is asked to return the $i$-th digit of a long integer, performance drops significantly as the length of the number increases. This suggests that current LLMs lack a consistent ability to simply find a digit. Note that the performance is good in the shorter S-range, which indicates that the models can at least comprehend the task instruction. In the XL-range, GPT-4o achieves only 20% accuracy, barely above the random guessing 10% baseline (since the correct answer is always a digit between 0 and 9). This fundamental limitation may explain why current LLMs struggle with numerical understanding and processing, especially as task complexity and input length increase. If a model cannot reliably identify a specific digit in a given number, it casts doubt on its ability to generalize to more complex arithmetic tasks, such as addition.
@@
We also have some interesting observations: (1) LLMs find the “max-hard” task easier than “max” with integer inputs. The difference between the tasks is that in the max task, the two numbers often differ in length, whereas in max-hard, they are always the same length and share some left-most digits, requiring more digits to be compared. While max-hard intuitively seems more difficult, models actually perform better on it. This is likely because they struggle to effectively use sequence length information, as reflected in their weaker performance on the “length” tasks in the longer ranges. It suggests that models might process tasks in different ways from humans. They could have to compare two numbers digit by digit. In this situation, the “harder” subtasks are actually easier because the numbers are already aligned. (2) GPT-4o and GPT-4o-mini show nearly identical performance across most tasks, similar to the comparison between Qwen2-72B and Qwen2-7B. This suggests that once a model reaches a certain size, NUPA performance relies more on factors like architecture, training strategies, data diversity, and post-training refinements, rather than simply on increasing model size.